A spring with spring constant of 175 N/m is used to pull a 10 kg crate across a level floor. If the coefficient of friction between the crate and the ground is .3 and the crate gets an acceleration of 4.5 m/s2, how far does the spring get stretched?
1
Expert's answer
2020-11-25T10:50:46-0500

Let's apply the Newton's Second Law of Motion:


"\sum F_x=ma_x,""F-F_{fr}=ma,""kx-\mu mg=ma."

From this equation we can find the elongation of the string:


"x=\dfrac{m(\mu g+a)}{k},""x=\dfrac{10\ kg\cdot(0.3\cdot 9.8\ \dfrac{m}{s^2}+4.5\ \dfrac{m}{s^2})}{175\ \dfrac{N}{m}}=0.42\ m."

Answer:

"x=0.42\ m."